OpenStudy (theedwardsfamily):

Which answer best describes the solution to the equation 5z + 4 = 4 + 5z? A. one solution B. no solution C. an infinite number of solutions D. z = 20

OpenStudy (theedwardsfamily):

@pooja195 @TheSmartOne

pooja195 (pooja195):

Which do you think?

OpenStudy (anonymous):

I think it is A. Try z=2.

OpenStudy (theedwardsfamily):

b?

OpenStudy (theedwardsfamily):

irdk

pooja195 (pooja195):

thats correct @gigi2002

OpenStudy (theedwardsfamily):

why is it correct?

OpenStudy (theedwardsfamily):

@pooja195

OpenStudy (anonymous):

when you do z=2 it makes since with the equation. which means it would be A. only one solution

OpenStudy (theedwardsfamily):

ok. thnx

OpenStudy (anonymous):

your welcome any time

TheSmartOne (thesmartone):

Its not A or B...

TheSmartOne (thesmartone):

5z + 4 = 4 + 5z Solve for Z.

OpenStudy (anonymous):

what proof do you have the it is not a or b

pooja195 (pooja195):

nvm it C there can be many solutions because when you slove it you can have any number in place

OpenStudy (theedwardsfamily):

okkk

TheSmartOne (thesmartone):

@gigi2002 You can plug in any number from -\(\infty\) till +\(\infty\) and the question will still be correct.

pooja195 (pooja195):

Yay!! I was right xD

TheSmartOne (thesmartone):

^ Well before you said B was correct :P

OpenStudy (anonymous):

oh so i guess i got it wrong

TheSmartOne (thesmartone):

5z + 4 = 4 + 5z You see? You can subtract 4 from both sides and get 5z = 5z And so basically any number can satisfy that.

pooja195 (pooja195):

my mistake sorry :(

TheSmartOne (thesmartone):

Its ok :) Everyone makes mistakes :D

OpenStudy (theedwardsfamily):

So the answr would be C right?

OpenStudy (anonymous):

yes

OpenStudy (theedwardsfamily):

ok